# CyC3D: Fine-grained Controllable 3D Generation via Cycle Consistency Regularization

### v1

@@ Abstract
Despite the remarkable progress of 3D generation, achieving controllability, i.e., ensuring consistency between generated 3D content and input conditions like edge and depth, remains a significant challenge. Existing methods often struggle to maintain accurate alignment, leading to noticeable discrepancies. To address this issue, we propose CyC3D, a new framework that enhances controllable 3D generation by explicitly encouraging cyclic consistency between the second-order 3D content, generated based on extracted signals from the first-order generation, and its original input controls. Specifically, we employ an efficient feed-forward backbone that can generate a 3D object from an input condition and a text prompt. Given an initial viewpoint and a control signal, a novel view is rendered from the generated 3D content, from which the extracted condition is used to regenerate the 3D content. This re-generated output is then rendered back to the initial viewpoint, followed by another round of control signal extraction, forming a cyclic process with two consistency constraints. View consistency ensures coherence between the two generated 3D objects, measured by semantic similarity to accommodate generative diversity. Condition consistency aligns the final extracted signal with the original input control, preserving structural or geometric details throughout the process. Extensive experiments on popular benchmarks demonstrate that CyC3D significantly improves controllability, especially for fine-grained details, outperforming existing methods across various conditions (e.g., +14.17% PSNR for edge, +6.26% PSNR for sketch).
@@ 1 Introduction
Figure 1: Visualization of input and reconstructed conditions. We feed the input condition to the generation model first and use the generated 3D contents to render the image at the input viewpoint to extract the reconstructed conditions. The difference between the input and the reconstructed conditions can reveal whether the fine-grained controls are reserved in the results.
Figure 2: Brief overview of our proposed CyC3D. We first feed the condition on the original reference view to the generation model and extract the conditions from the rendered results on randomly sampled novel views. The extracted conditions are then fed to the generation model again and rendered back to the original reference view. The extracted condition of the rendered results should be consistent with the input condition on the reference view.
The creation of 3D models from text descriptions (text-to-3D) or image collections (image-to-3D) is a fundamental task in computer graphics, attracting growing interest from a wide range of domains, including VR/AR [2], digital game design [32], and filmmaking [30]. The mainstream approaches can be broadly categorized into two groups: optimization-based methods and feed-forward methods. Optimization-based 3D generation methods leverage Score Distillation Sampling (SDS) [20] to iteratively refine 3D objects based on textual or visual inputs, effectively distilling the rich prior knowledge embedded in large pre-trained diffusion models [24]. In contrast, feed-forward 3D generation approaches [19, 28, 11, 12, 36, 29, 5] generate 3D content in a single inference pass, significantly improving efficiency and making real-time applications more feasible.
@@
To advance research in controllable 3D generation, MVControl [15] extends ControlNet, originally designed for conditioned 2D image generation, to a multi-view diffusion model [25]. It first predicts four-view images, which are then processed by a multi-view Gaussian reconstruction model [27] to generate the final 3D representation. In contrast, ControLRM [31] aims to develop an end-to-end controllable 3D generation framework by replacing the original image encoder of a large reconstruction model (LRM) with a conditional encoder, enabling more precise conditioning. While these studies have demonstrated a degree of controllability, a significant challenge remains, i.e., achieving precise and fine-grained conditional control. As illustrated in Fig. 1, the extracted signals from the 3D models generated by MVControl and ControLRM exhibit noticeable discrepancies from the original input controls, underscoring the need for more refined and accurate controllable 3D generation.
@@
To address this challenge, we propose CyC3D, a novel and general training framework that enhances the controllability of 3D generation. Our approach extends a feed-forward 3D generation network into a cyclic paradigm, enabling more precise and consistent adherence to input control signals, as illustrated in Fig. 2. Specifically, given an input control condition on a reference view, the generated 3D content is first rendered from randomly sampled new viewpoints. From these rendered images, the corresponding control conditions are extracted and fed back into the generation model as inputs, ensuring that the model learns to maintain control consistency across views. Subsequently, the model regenerates the 3D content, which is then rendered back to the initial reference viewpoint, where the extracted control conditions should match the original input condition. To further refine controllability, we introduce two complementary consistency constraints: View Cycle Consistency, which ensures that the two generated contents have similar semantics in embedded space since they are generated with the same prompt and highly coupled conditions, and Condition Cycle Consistency, which enforces strict alignment between the extracted control signals from the final rendered images and the original input condition. We compare our CyC3D against state-of-the-art approaches on GSO and ABO datasets [31] for zero-shot evaluation. Experimental results demonstrate that our framework significantly enhances controllability across all benchmarks while preserving competitive generation quality compared to existing methods.
We highlight the contributions of this work as follows:
@@
  • New Insight and Framework: We reveal that existing efforts in controllable 3D generation still perform poorly on fine-grained controllability, and propose a novel framework (CyC3D) to handle the problem, which extends the original single-stage novel view synthesis pipeline to a double-stage one with cycle consistency (Fig. 2).
  • Cyclic Consistency Feedback: To refine controllability, we propose two complementary consistency constraints: condition cycle consistency and view cycle consistency. The former regularizes the consistency between input and reconstructed conditions, the latter enhances the semantic consistency of different stages.
  • Evaluation and Promising Results: We provide a comprehensive zero-shot evaluation of the controllability on GSO/ABO benchmarks, and demonstrate that CyC3D comprehensively outperforms existing methods.
@@ 2 Related Work
Text-/Image-to-3D Generation. 3D generation can be broadly categorized into optimization-based and feed-forward approaches: (1) Optimization-based methods offer a viable alternative to traditional techniques by circumventing the need for extensive text-to-3D datasets. DreamFusion [20] exemplifies this approach by utilizing SDS loss to optimize neural fields with diffusion priors, resulting in high-quality 3D assets. Despite their effectiveness, these methods often encounter the Janus problem, which arises from inconsistencies between 2D supervision and 3D geometry. (2) Feed-forward models have emerged as a prominent approach for efficient 3D generation, leveraging extensive 3D datasets. Techniques such as LRM [12] and TripoSR [29] utilize transformer architectures to predict NeRF from single-view images, demonstrating high generalization capabilities. Multi-view methods extend these approaches by synthesizing higher-quality 3D models from single-view images using multi-view diffusion models, as demonstrated by SyncDreamer [16] and MVDream [25].
Controllable 3D Generation. Achieving controllability in 3D generation remains a critical objective in the field. Techniques developed for controllable 2D generation have provided valuable insights and methods that are applicable to the more complex domain of controllable 3D generation. Inspired by ControlNet [35], MVControl [15] integrates a trainable control network with a multi-view diffusion model to facilitate controllable multi-view image generation. This method combines coarse Gaussian generation with SDS-based refinement but faces challenges related to model generalization and procedural complexity. ControLRM [31] addresses these limitations with its end-to-end feed-forward model, ensuring rapid inference and enhanced control precision, thus offering an efficient and effective solution for controllable 3D generation. Building on the advances of ControlNet++ [14] and ControLRM [35], we improve conditional controls to achieve more sophisticated and nuanced control over the generated 3D assets.
3 CyC3D
Figure 3: Overall framework of our proposed CyC3D. (a) summarize the procedure of our CyC3D. (b) shows the condition extractor utilized for different conditional controls (Sec. 3.2). (c) illustrates the condition cycle consistency regularization (Sec. 3.3). (d) introduces the proposed view cycle consistency regularization (Sec. 3.4).
As shown in Fig. 3, we present the overall framework of our proposed CyC3D. In Sec. 3.1, we first introduce the feed-forward backbone network for controllable 3D generation in our CyC3D. Then we introduce the condition extractor (Fig. 3(b)) for different conditional controls (e.g. edge, sketch, depth, normal) in Sec. 3.2. In Sec. 3.3, we introduce the design of condition cycle consistency regularization (Fig. 3(c)) to enhance the fined-grained controls. In Sec. 3.4, we further introduce the proposed view cycle consistency regularization (Fig. 3(d)) to stabilize the training and avoid overfitting on the input condition. Finally, we introduce the overall loss in Sec. 3.5.
@@ 3.1 Generation Backbone
For controllable 3D generation, our CyC3D can be combined with arbitrary feed-forward backbones (e.g. ControLRM [31], MVControl [15]). In default, we select ControLRM [31] as our backbone model. ControLRM contains a 2D condition extractor and a 3D triplane decoder. (1) The conditional generator in ControLRM is designed to take both text and 2D visual conditions as inputs. This generator can utilize either a transformer or diffusion backbone to create initial 2D latent representations, which are subsequently transformed into 3D models. (2) The 3D triplane decoder is inherited from a pre-trained large reconstruction model [12]. The triplane transformer decoder employs a unique representation of 3D data by decomposing it into three orthogonal planes (X, Y, and Z). Each plane is processed using transformer layers, which utilize self-attention mechanisms to capture spatial relationships and dependencies within the data.
3.2 Condition Extractor
To ensure the gradient back-propagation of cycle consistency in our CyC3D framework during training, the utilized condition extractor in Fig. 2 should be differentiable. The design of the condition extractor for each condition control is shown as follows:
Canny Condition. As shown in the figure, we extract edge maps using the Canny algorithm [4]. Following [14], a differentiable Canny extractor can implemented based on Kornia [23] to extract edge maps.
Sketch Condition. Following ControlNet [35], we utilize a pre-trained CNN network to extract sketch conditions from images. Since the CNN network is differentiable, the gradients can be passed backward through the network and back-propagated to the remaining modules.
Depth Condition. In analogy with the sketch condition, we also utilize a pre-trained foundation model (Depth Anything [33]) to extract the depth condition from rendered images. The whole process is differentiable in the cycle consistency pipeline.
Normal Condition. In practice, we find that the existing normal estimation models [1] utilized in ControlNet [35] tends to predict over-smoothed normal maps. It may result in a large gap between the exact 3D shape of the rendered multi-view images and the estimated normal maps. Consequently, we adopt an alternative solution to predict the depth map first and then convert the depth map to normal map via a differentiable module [13]. The great zero-shot generalization ability of large depth-estimation models like Depth-Anything [33] can effectively capture the 3D prior from the image and thus generate accurate normal map.
@@ 3.3 Condition Cycle Consistency Regularization
Given a 3D point $x\in\mathbb{R}^{3}$, the color and density filed $(c,\sigma)$ can be obtained via sampling the corresponding features on the triplane $P_{i}$: $(c,\sigma)=f_{\text{mlp}}(P_{i},x)$. By sampling the points along the viewpoint $T_{r}\in\mathbb{R}^{4\times 4}$, we can render the image on view $T_{r}$ via volume rendering [17]:
$\small\hat{I_{r}},\hat{D_{r}}=f_{\text{render}}(P_{i},T_{r}),$ (2)
@@
Figure 4: Visualization results of our CyC3D. The input conditions and the rendered images at different views are visualized. Our CyC3D can generate high quality results with consistent conditional controls.
@@
$\small\overline{I_{i}},\overline{D_{i}}=f_{\text{render}}(P_{r},T_{i}),$ (5)
where $\overline{I_{i}}$ is the rendered image on the input reference view, and $\overline{D_{i}}$ is the rendered depth map. Then we can extract the conditon map from the rendered image $\overline{I_{i}}$:
$\small\overline{C_{i}}=f_{\text{cond}}(\overline{I_{i}}),$ (6)
@@
As shown in Fig. 2 and 3, we can calculate the difference between the reconstructed condition and the input condition to regularize the fine-grained controls. Depending on the type of conditional controls, we also propose two different kinds of conditional cycle-consistency feedbacks as follows:
2D Conditional Control Feedback: If we only consider 2D conditional controls extracted from the rendered images (e.g. Canny map extracted from RGB image), we can use a condition extractor to reconstruct the control signal. The generated condition map $\overline{C_{i}}$ via Eq. 6 should be consistent with the original condition map $C_{v}$ on the input reference view.
$\small L_{\text{cond}}=\sum_{T_{r}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}.$ (7)
If the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, the two-step inference in cycle consistency will be degraded to a single one (as shown in Fig. 3 (c)), saving the computation cost:
$\small\hat{I_{i}},\hat{D_{i}}=f_{\text{render}}(f_{\text{gen}}(C_{v},C_{t},T_{i}),T_{i}),$ (8)
$\small\hat{C_{i}}=f_{\text{cond}}(\hat{I_{i}}),$ (9)
@@
$\small L_{\text{cond}}=\sum_{T_{r}\neq T_{i}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}+\|\hat{C_{i}}-C_{v}\|_{2}^{2}.$ (10)
Figure 5: Qualitative comparison among our CyC3D and the other state-of-the-art controllable 3D generation methods.
3D Conditional Control Feedback: If we consider 3D conditional controls like depth or normal conditions, the 3D shape of the generated objects should be regularized as well for a fine-grained control.
@@
$\small L_{\text{cond-d}}=\sum_{T_{r}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2},$ (11)
where $f_{\text{norm}}$ is the depth normalizing function. Since the rendered depth $\overline{D_{i}}$ and the input depth $C_{d}$ have a gap in the scale of depth, we utilize $f_{\text{norm}}$ to normalize them to the same scale and calculate a scale-agnostic loss following [22].
In analogy with Eq. 10, when the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation. We can modify Eq. 11 as follows:
$\small L_{\text{cond-d}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2}+\|f_{\text{norm}}(\hat{D_{i}})-C_{d}\|_{2}^{2},$ (12)
@@
(2) If the input condition is normal map ($C_{v}=C_{n}$), the explicit depth-normal consistency can be further calculated as follows:
$\small L_{\text{cond-n}}=\sum_{T_{r}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2},$ (13)
where $f_{\text{d2n}}$ is a differentiable module [13] that can converts the depth map to a unified normal map.
In analogy with Eq. 10 and 12, when the randomly sampled camera pose $T_{r}$ equals to the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation. We can thus modify Eq. 13 as follows:
$\small L_{\text{cond-n}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2}+\|f_{\text{d2n}}(\hat{D_{i}})-C_{n}\|_{2}^{2},$ (14)

### v2

@@ Abstract
Despite the remarkable progress of 3D generation, achieving controllability, i.e., ensuring consistency between generated 3D content and input conditions like edge and depth, remains a significant challenge. Existing approaches often struggle to maintain accurate alignment, leading to noticeable discrepancies. To address this issue, we propose CyC3D, a new framework designed to enhance controllable 3D generation by explicitly encouraging cyclic consistency during training between the second-order 3D content, generated based on extracted signals from the first-order generation, and its original input controls. Specifically, we employ an efficient feed-forward backbone that can generate a 3D object from an input condition and a text prompt. Given an initial viewpoint and a control signal, a novel view is rendered from the generated 3D content, from which the extracted condition is used to regenerate the 3D content. This re-generated output is then rendered back to the initial viewpoint, followed by another round of control signal extraction, forming a cyclic process with two consistency constraints. View consistency ensures coherence between the two generated 3D objects, measured by semantic similarity to accommodate generative diversity. Condition consistency aligns the final extracted signal with the original input control, preserving structural or geometric details throughout the process. Extensive experiments on zero-shot GSO/ABO benchmarks demonstrate that CyC3D significantly improves controllability, especially for fine-grained details, outperforming existing methods across various conditions (e.g., +14.17% PSNR for edge, +6.26% PSNR for sketch).
@@ 1 Introduction
Figure 1: Brief overview of our proposed CyC3D.
The creation of 3D models from text descriptions (text-to-3D) or image collections (image-to-3D) is a fundamental task in computer graphics, attracting growing interest from a wide range of domains, including VR/AR (Behravan 2025), digital game design (Xu et al. 2024b), and filmmaking (Wang et al. 2025). The mainstream approaches can be broadly categorized into two groups: optimization-based methods and feed-forward methods. Optimization-based 3D generation methods leverage Score Distillation Sampling (SDS) (Poole et al. 2022) to iteratively refine 3D objects based on textual or visual inputs, effectively distilling the rich prior knowledge embedded in large pre-trained diffusion models (Rombach et al. 2022). In contrast, feed-forward 3D generation approaches (Nichol et al. 2022; Tang et al. 2023b; Hong et al. 2024, 2023; Zou et al. 2024; Tochilkin et al. 2024; Chen et al. 2025) generate 3D content in a single inference pass, significantly improving efficiency and making real-time applications more feasible.
@@
To advance research in controllable 3D generation, MVControl (Li et al. 2024) extends ControlNet, originally designed for conditioned 2D image generation, to a multi-view diffusion model (Shi et al. 2023). It first predicts four-view images, which are then processed by a multi-view Gaussian reconstruction model (Tang et al. 2025a) to generate the final 3D representation. In contrast, ControLRM (Xu et al. 2024a) aims to develop an end-to-end controllable 3D generation framework by replacing the original image encoder of a large reconstruction model (LRM) (Hong et al. 2023) with a conditional encoder, enabling more precise conditioning. The inference pipeline of these controllable 3D generation works is summarized in Fig. 1 (a). While these studies have demonstrated a degree of controllability, a significant challenge remains, i.e., , achieving precise and fine-grained conditional control. As illustrated in Fig. 1 (b), the extracted signals from the 3D models generated by MVControl and ControLRM exhibit noticeable discrepancies from the original input controls, underscoring the need for more refined and accurate controllable 3D generation.
@@
To address this challenge, we propose CyC3D, a novel and general training framework that enhances the controllability of 3D generation. Our approach extends a feed-forward 3D generation network into a cyclic paradigm, enabling more precise and consistent adherence to input control signals, as illustrated in Fig. 1 (c). Specifically, given an input control condition on a reference view, the generated 3D content is first rendered from randomly sampled new viewpoints. From these rendered images, the corresponding control conditions are extracted and fed back into the generation model as inputs, ensuring that the model learns to maintain control consistency across views. Subsequently, the model regenerates the 3D content, which is then rendered back to the initial reference viewpoint, where the extracted control conditions should match the original input condition. To further refine controllability, we introduce two complementary consistency constraints: View Cycle Consistency, which ensures that the two generated contents have similar semantics in embedded space since they are generated with the same prompt and highly coupled conditions, and Condition Cycle Consistency, which enforces strict alignment between the extracted control signals from the final rendered images and the original input condition. We compare our CyC3D against state-of-the-art approaches on GSO and ABO datasets (Xu et al. 2024a) for zero-shot evaluation. Experimental results demonstrate that our framework significantly enhances controllability across all benchmarks while preserving competitive generation quality compared to existing methods.
Figure 2: Training framework of our proposed CyC3D. It contains two kinds of regularization: condition cycle consistency regularization (Sec. 3.3) and view cycle consistency regularization (Sec. 3.4).
@@
We highlight the contributions of this work as follows: (1) New Insight and Framework: We reveal that existing efforts in controllable 3D generation still perform poorly on fine-grained controllability, and propose a novel framework (CyC3D) to handle the problem, which extends the original single-stage novel view synthesis pipeline to a double-stage one with cycle consistency. (2) Cyclic Consistency Feedback: To refine controllability, we propose two complementary consistency constraints: condition cycle consistency and view cycle consistency. The former regularizes the consistency between input and reconstructed conditions, the latter enhances the semantic consistency of different stages. (3) Evaluation and Promising Results: We provide a comprehensive zero-shot evaluation of the controllability on GSO/ABO benchmarks, and demonstrate that CyC3D comprehensively outperforms existing methods.
@@ 2 Related Work
Text-/Image-to-3D Generation. 3D generation can be broadly categorized into optimization-based and feed-forward approaches: (1) Optimization-based methods offer a viable alternative to traditional techniques by circumventing the need for extensive text-to-3D datasets. DreamFusion (Poole et al. 2022) exemplifies this approach by utilizing SDS loss to optimize neural fields with diffusion priors, resulting in high-quality 3D assets. Despite their effectiveness, these methods often encounter the Janus problem, which arises from inconsistencies between 2D supervision and 3D geometry. (2) Feed-forward models have emerged as a prominent approach for efficient 3D generation, leveraging extensive 3D datasets. Techniques such as LRM (Hong et al. 2023) and TripoSR (Tochilkin et al. 2024) utilize transformer architectures to predict NeRF from single-view images, demonstrating high generalization capabilities. Multi-view methods extend these approaches by synthesizing higher-quality 3D models from single-view images using multi-view diffusion models, as demonstrated by SyncDreamer (Liu et al. 2023) and MVDream (Shi et al. 2023). Though Cycle3D (Tang et al. 2025b) builds a cycle-consistent pipeline of diffusion and reconstruction models, the whole pipeline is built in the inference model. Whereas our cycle-consistency is used for training a real-time feed-forward generation model.
Controllable 3D Generation. Achieving controllability in 3D generation remains a critical objective in the field. Techniques developed for controllable 2D generation have provided valuable insights and methods that are applicable to the more complex domain of controllable 3D generation. Inspired by ControlNet (Zhang, Rao, and Agrawala 2023), MVControl (Li et al. 2024) integrates a trainable control network with a multi-view diffusion model to facilitate controllable multi-view image generation. This method combines coarse Gaussian generation with SDS-based refinement but faces challenges related to model generalization and procedural complexity. ControLRM (Xu et al. 2024a) addresses these limitations with its end-to-end feed-forward model, ensuring rapid inference and enhanced control precision, thus offering an efficient and effective solution for controllable 3D generation. Building on the advances of ControlNet++ (Li et al. 2025) and ControLRM (Zhang, Rao, and Agrawala 2023), we improve conditional controls to achieve more sophisticated and nuanced control over the generated 3D assets.
3 Method
@@ 3.1 Generation Backbone
For controllable 3D generation, our CyC3D can be combined with arbitrary feed-forward backbones (e.g. ControLRM (Xu et al. 2024a), MVControl (Li et al. 2024)). In default, we select ControLRM (Xu et al. 2024a) as our backbone model. ControLRM contains a 2D condition extractor and a 3D triplane decoder. The conditional generator in ControLRM is designed to take both text and 2D visual conditions as inputs. This generator can utilize either a transformer or diffusion backbone to create initial 2D latent representations, which are subsequently transformed into 3D models. The 3D triplane decoder is inherited from a pre-trained large reconstruction model (Hong et al. 2023). The triplane transformer decoder employs a unique representation of 3D data by decomposing it into three orthogonal planes (X, Y, and Z). Each plane is processed using transformer layers, which utilize self-attention mechanisms to capture spatial relationships and dependencies within the data.
Figure 3: Visualization results of our CyC3D. The input conditions and the rendered images at different views are visualized. Our CyC3D can generate high-quality results with consistent conditional controls.
3.2 Condition Extraction
To ensure the gradient back-propagation of cycle consistency in our CyC3D framework during training, the utilized condition extractor should be fully differentiable. The design of the condition extractor for each condition control is shown as follows:
Canny Condition. We extract edge maps using the Canny algorithm (Canny 1986). Following (Li et al. 2025), a differentiable Canny extractor can be implemented based on Kornia (Riba et al. 2020) to extract edge maps.
Sketch Condition. Following ControlNet (Zhang, Rao, and Agrawala 2023), we utilize a pre-trained CNN network to extract sketch conditions from images. Since the CNN network is differentiable, the gradients can be passed backward through the network.
Depth Condition. In analogy with the sketch condition, we also utilize a pre-trained foundation model (Depth Anything (Yang et al. 2024)) to extract the depth condition from rendered images. As an alternative, we can also normalize the rendered depth map via volume rendering to extract the geometric prior from the 3D content.
Normal Condition. In practice, we find that the existing normal estimation models (Bae, Budvytis, and Cipolla 2021) utilized in ControlNet (Zhang, Rao, and Agrawala 2023) tend to predict over-smoothed normal maps. It may result in a large gap between the exact 3D shape of the rendered multi-view images and the estimated normal maps. Consequently, we adopt an alternative solution to predict the depth map first and then convert the depth map to a normal map via a differentiable module (Huang et al. 2021). As an alternative, we also extract the normal from the rendered depth maps via volume rendering.
@@ 3.3 Condition Cycle Consistency Regularization
Given a 3D point $x\in\mathbb{R}^{3}$, the color and density filed $(c,\sigma)$ can be obtained via sampling the corresponding features on the triplane $P_{i}$: $(c,\sigma)=f_{\text{mlp}}(P_{i},x)$. By sampling the points along the viewpoint $T_{r}\in\mathbb{R}^{4\times 4}$, we can render the image on view $T_{r}$ via volume rendering (Mildenhall et al. 2021):
$\hat{I_{r}},\hat{D_{r}}=f_{\text{render}}(P_{i},T_{r}),$ (2)
@@
$\overline{I_{i}},\overline{D_{i}}=f_{\text{render}}(P_{r},T_{i}),$ (5)
where $\overline{I_{i}}$ is the rendered image on the input reference view, and $\overline{D_{i}}$ is the rendered depth map. Then we can extract the condition map from the rendered image $\overline{I_{i}}$:
$\overline{C_{i}}=f_{\text{cond}}(\overline{I_{i}}),$ (6)
@@
As shown in Fig. 2, we can calculate the difference between the reconstructed condition and the input condition to regularize the fine-grained controls. Depending on the type of conditional controls, we also propose two different kinds of conditional cycle-consistency feedback as follows:
2D Conditional Control Feedback. If we only consider 2D conditional controls extracted from the rendered images (e.g., Canny map extracted from RGB image), we can use a condition extractor to reconstruct the control signal. The generated condition map $\overline{C_{i}}$ via Eq. 6 should be consistent with the original condition map $C_{v}$ on the input reference view.
$L_{\text{cond}}=\sum_{T_{r}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}.$ (7)
If the randomly sampled camera pose $T_{r}$ equals the input viewpoint $T_{i}$, the two-step inference in cycle consistency will be degraded to a single one (as shown in Fig. 2 (c)), saving the computation cost:
$\hat{I_{i}},\hat{D_{i}}=f_{\text{render}}(f_{\text{gen}}(C_{v},C_{t},T_{i}),T_{i}),$ (8)
[…]
@@
$L_{\text{cond}}=\sum_{T_{r}\neq T_{i}}\|\overline{C_{i}}-C_{v}\|_{2}^{2}+\|\hat{C_{i}}-C_{v}\|_{2}^{2}.$ (10)
Figure 4: Qualitative comparison among our CyC3D and the other state-of-the-art controllable 3D generation methods. In the 1st row, the baseline methods fail to generate consistent content with the text prompts. In the 2nd row, ControLRM generates a fainter color compared with our CyC3D.
3D Conditional Control Feedback. If we consider 3D conditional controls, such as depth or normal conditions, the 3D shape of the generated objects should also be regularized for fine-grained control.
@@
$L_{\text{cond-d}}=\sum_{T_{r}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2},$ (11)
where $f_{\text{norm}}$ is the depth normalizing function. Since the rendered depth $\overline{D_{i}}$ and the input depth $C_{d}$ have a gap in the scale of depth, we utilize $f_{\text{norm}}$ to normalize them to the same scale and calculate a scale-agnostic loss following (Ranftl et al. 2020).
In analogy with Eq. 10, when the randomly sampled camera pose $T_{r}$ equals the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation. We can modify Eq. 11 as follows:
$L_{\text{cond-d}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{norm}}(\overline{D_{i}})-C_{d}\|_{2}^{2}+\|f_{\text{norm}}(\hat{D_{i}})-C_{d}\|_{2}^{2},$ (12)
@@
(2) If the input condition is a normal map ($C_{v}=C_{n}$), the explicit depth-normal consistency can be further calculated as follows:
$L_{\text{cond-n}}=\sum_{T_{r}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2},$ (13)
where $f_{\text{d2n}}$ is a differentiable module (Huang et al. 2021) that can convert the depth map to a unified normal map.
In analogy with Eq. 10 and 12, when the randomly sampled camera pose $T_{r}$ equals the input viewpoint $T_{i}$, there is no need to run the inference twice during propagation. We can thus modify Eq. 13 as follows:
$L_{\text{cond-n}}=\sum_{T_{r}\neq T_{i}}\|f_{\text{d2n}}(\overline{D_{i}})-C_{n}\|_{2}^{2}+\|f_{\text{d2n}}(\hat{D_{i}})-C_{n}\|_{2}^{2},$ (14)
@@
Methods Edge Sketch Depth Normal
PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
GSGEN (Chen et al. 2024) 13.01 0.780 0.0526 15.07 0.772 0.0322 0.1277 0.1268 0.0311 0.0230 0.0497
GaussianDreamer (Yi et al. 2024) 12.41 0.766 0.0597 14.58 0.792 0.0362 0.0740 0.0947 0.0391 0.0225 0.0366
DreamGaussian (Tang et al. 2023a) 9.94 0.686 0.1048 14.42 0.765 0.0380 0.0980 0.0916 0.0406 0.0216 0.0335
VolumeDiffusion (Tang et al. 2023b) 13.25 0.818 0.0497 16.50 0.836 0.0231 0.1202 0.0953 0.0531 0.0190 0.0172
3DTopia (Hong et al. 2024) 9.63 0.704 0.1134 15.54 0.801 0.0287 0.1193 0.1064 0.0420 0.0268 0.0372
MVControl (Li et al. 2024) 10.44 0.723 0.0952 14.51 0.746 0.0370 0.0555 0.0645 0.0160 0.0209 0.0315
ControLRM-T (Xu et al. 2024a) 12.43 0.836 0.0607 17.17 0.836 0.0202 0.0567 0.0713 0.0599 0.0120 0.0359
ControLRM-D (Xu et al. 2024a) 12.47 0.837 0.0600 17.23 0.837 0.0197 0.0383 0.0546 0.0650 0.0119 0.0357
Table 1: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on GSO benchmark. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. The best results are highlighted in bold. We provide the results of four different conditions: edge, sketch, depth, and normal.
Methods Edge Sketch Depth Normal
PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
GSGEN (Chen et al. 2024) 12.75 0.737 0.0614 14.25 0.744 0.0384 0.1004 0.0967 0.0463 0.0171 0.0479
GaussianDreamer (Yi et al. 2024) 11.92 0.736 0.0699 13.75 0.734 0.0439 0.0809 0.1061 0.0521 0.0188 0.0463
DreamGaussian (Tang et al. 2023a) 9.38 0.683 0.1178 12.90 0.716 0.0523 0.1267 0.0762 0.0433 0.0188 0.0349
VolumeDiffusion (Tang et al. 2023b) 12.58 0.734 0.0619 15.03 0.780 0.0321 0.1007 0.1198 0.0671 0.0188 0.0264
3DTopia (Hong et al. 2024) 8.48 0.633 0.1469 14.41 0.741 0.0372 0.1916 0.1844 0.0559 0.0346 0.0418
MVControl (Li et al. 2024) 9.52 0.656 0.1161 13.34 0.686 0.0479 0.0454 0.0467 0.0181 0.0133 0.0350
ControLRM-T (Xu et al. 2024a) 10.49 0.552 0.0909 15.77 0.777 0.0272 0.0514 0.0497 0.0700 0.0136 0.0572
ControLRM-D (Xu et al. 2024a) 11.51 0.592 0.0725 15.80 0.778 0.0269 0.0493 0.0545 0.0661 0.0158 0.0548
Table 2: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on ABO benchmark. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. The best results are highlighted in bold. We provide the results of four different conditions: edge, sketch, depth, and normal.